Perth Artifactory

Subcommittee:laser-minutes-20161214

Laser subcommittee meeting 2016-12-21
  • Issues since previously, or continuing
    • LC1290
      • Bed level check
        • 2016-12-05: LEM+JM checked: bed is level
          • Slats supporting the bed have MDF char and move around
          • New honeycomb?
      • but: Mirror alignment needed at 2016-12-05: masking tape shows the beam seeming to go quite high into the aperture at the top of the head at all areas of the travel
      • TODO: Check chiller
        • water level
        • Does temperature stay on target or creep up?
        • temp snapshot: 19.2C, 18.8C
      • New tube test install
        • Pick a downtime day: Wed 20161228-0900 ; with a lunchtime deadline for rollback
    • LG500
      • temp snapshot: 16.0C, 16.4C
    • Can we find/replace the missing access keys? Only 2 of 6? remain
    • Monitor/backup PC config
    • No fresh spare lens: want 2 * 50.8mm + 2 * 63.5mm
      • Regarded as a consumable
      • Want list of spares on the wiki
      • Resolved: buy some!
    • Open up artifactory-laser list view topics: agreed; DONE